OpenStudy (anonymous):

Find the variation constant and an equation of the variation where y varies inversely as x and y=2 when x=20.

OpenStudy (amistre64):

division is inversion here: y = k/x 2 = k/20 k = 40 y = 40/x

OpenStudy (anonymous):

I am confused what the variation constant is .

OpenStudy (amistre64):

the constant of variation is the term that is introduce into the equation to provide some sense of stability to it. In this instance; that constant = 40; how else can we establish the relation between x and y?

OpenStudy (anonymous):

Thank you very much

OpenStudy (anonymous):

x and y are parallel ?

OpenStudy (amistre64):

As y increases; x has to decrease; its in inverted relationship; The only way to do that in math is to either subtract of divide; and this implies division :) y = k/x establishes the relationship we know that when y = 2; x=20, we need another term in the equation to make that statment true; that term is the constant of variation :)

OpenStudy (anonymous):

note to self - never do algebra with an online school, :) Thanks again

OpenStudy (amistre64):

when x = 1; y = 40 when x = 2; y = 20 when x = 4; y = 10 you notice here that since k=40, we have increasing x values and decreasing y values; they have an inverse relationship now :) Yeah, online school might sound convenient at times, but I tried it once and dint like it

OpenStudy (anonymous):

I think i get it, so when x goes up y goes down ... like a pulley so y=60 would be x=1 if x=6 when y=10?

OpenStudy (amistre64):

to use the equation we determined: y = 40/x if y = 60; then x = 60/40 = 6/4 = 3/2

OpenStudy (amistre64):

but yes; like a pulley :)